3s squared +2 = -62
I subtracted 2 so that 3s squared = -60 and divided -60 by 3, giving me -20. I then had s squared = -20 and, since you can't have negative square roots, i put S= + or - 2 root 5, but im not sure if that is right.

Answer by edjones(7311)

3s^2+2=-62
3s^2=-64
s^2=-64/3
s=+-8sqrt(-1/3)
=8sqrt(1/3)i [i=sqrt(-1)]
sqrt(1/3)=1/sqrt(3)=1/sqrt(3) * sqrt(3)/sqrt(3) = sqrt(3)/3 no sqrt in denominator.
s=8sqrt(3)i/3, s=-8sqrt(3)i/3
.
Ed